OpenStudy (anonymous):

Solve the following system of equations. 3y = 12 2x + y = -16

OpenStudy (anonymous):

3y=12 2x+y=-16 we are gonna use substitution for the y and in order to do that we must isolate the y 2x+y=-16 -2x -2x ________________ y=-2x-16 Now that we've isolated the y we must substitute y into the top equation. 3(-2x-16)=12 Use the distributive property -6x-48=12 +48 +48 ____________-____ -6x= 60 divide negative 6 on both sides and you get x=-10<----- then you substitute x into any of the original equations so im gonna use the second equation 2(10)+y=-16 20+y=-16 -20 -20 _________________ y= -36<-----